# A Moreau Envelope Approach for LQR Meta-Policy Estimation

### v1

@@ I Introduction
Meta-learning facilitates quick adaptation to new tasks by applying previously acquired knowledge to new challenges [10, 11]. In RL, this evolves into Meta-RL, which trains agents to efficiently adapt to unknown environments [12, 13, 14, 15]. The Meta-RL framework aims to develop a meta-policy that allows for rapid learning across various tasks, fostering knowledge adaptation to new scenarios.
@@
Recently, in [19, 20, 21], the authors studied the meta-learning problem for LQR policy design. These works focused on meta-learning based on the MAML framework. In [21], the authors study MAML for the LQR problem in a single system and multi-task setting, which was later extended to a multi-system and multi-task setting by [19]. In their setup, data comes from a finite set of observed tasks, and it is used to design an LQR meta-policy that can be effectively and quickly adapted to unobserved tasks, but only local convergence is shown. Later, the authors in [20] studied system heterogeneity and provided global convergence guarantees.
When applying policy gradient techniques, MAML and ME define different surrogate cost functions and access different computational oracles. MAML uses the estimation of the gradient and Hessian of the cost function, whereas ME uses an approximate solution of an inner optimization problem via a first-order oracle. The inner loop accuracy parameter appears in the convergence guarantee, thereby providing the user with explicit control over the quality of the solution. Complied with this theoretical intuition, recent studies have shown that ME provides a better empirical performance than MAML in multi-task setups [16, 18, 17].
@@
The main contribution of this paper is a novel first-order meta-RL algorithm (MEMLQR) that computes LQR meta-policies that can be efficiently adapted for the control of uncertain linear systems. By integrating meta-learning principles, our approach improves policy adaptability, ensuring effective performance in various scenarios and quick convergence.
@@
    We propose a first-order iterative algorithm to optimize the defined augmented cost function in a client-server setup.
@@
    We show the convergence of the proposed algorithm to an approximate first-order stationary point. We also show that the policies generated by the algorithm will incur a finite cost for all the available system realizations (i.e., stability).
@@ II-A Preliminaries
such that,
$\mathcal{A}:=\mathopen{}\mathclose{{}\left(\mathcal{A}_{0}+\sum_{i=1}^{p}a_{i}\mathcal{A}_{i}}\right)\text{ and }\mathcal{B}:=\mathcal{B}_{0}+\sum_{j=1}^{q}b_{j}\mathcal{B}_{j},$ (2)
here, $x^{t}\in\mathbb{R}^{n}$ is the state and $u^{t}\in\mathbb{R}^{m}$ is the input to the system at time $t$, $a:=\mathopen{}\mathclose{{}\left(a_{1}\;a_{2}\;\cdots\;a_{p}}\right)\in\mathbb{R}^{p}$ and $b:=\mathopen{}\mathclose{{}\left(b_{1}\;b_{2}\;\cdots\;b_{q}}\right)\in\mathbb{R}^{q}$ are bounded uncertain parameters that determine the system dynamics in association with matrices $\mathcal{A}_{0}\in\mathbb{R}^{n\times n}$ and $\mathcal{B}_{0}\in\mathbb{R}^{n\times m}$, and sets of matrices, $\bar{\mathcal{A}}{}:=\mathopen{}\mathclose{{}\left(\mathcal{A}_{1},\mathcal{A}_{2},\cdots,\mathcal{A}_{p}}\right)\in\mathopen{}\mathclose{{}\left(\mathbb{R}^{n\times n}}\right)^{p}$ and ¯$\bar{\mathcal{B}}{}:=\mathopen{}\mathclose{{}\left(\mathcal{B}_{1},\mathcal{B}_{2},\cdots,\mathcal{B}_{q}}\right)\in\mathopen{}\mathclose{{}\left(\mathbb{R}^{n\times m}}\right)^{q}$. It is assumed that all possible realizations of the $\mathcal{A}-\mathcal{B}$ pair are controllable.
@@
subject to (3), where $Q\in\mathbb{R}^{n\times n}$ and $R\in\mathbb{R}^{m\times m}$ are the cost matrices such that $Q\succeq 0$ and $R\succ 0$. Without loss of generality, it is assumed that the cost matrices $Q$ and $R$ are the same for all the systems. This models a scenario where the controllers have the same goal or are trying to accomplish a similar task.
The optimal control policy for the above cost function is a linear feedback policy of form $u^{t}_{i}=-K_{i}x^{t}_{i}$ for a policy $K_{i}\in\mathbb{R}^{m\times n}$ [1, Chapter 3]. Therefore, the cost can be expressed as
@@ II-B Moreau Envelope-based LQR Meta-Policy Estimation
While the approach presented by the policy gradient iterations in (6) is efficient for the computation of approximate first-order optimal policies, here we focus on the task of learning good initialization that can be efficiently adapted for the LQR problem on unseen realizations of (3). We propose to use the Moreau Envelope (ME) approach for meta-learning [18, 17, 16] that defines an optimization problem as
@@
We propose a gradient-based iterative framework termed MEMLQR: Moreau Envelope based Meta Linear Quadratic Regulator, shown in Algorithm 1, to solve the bilevel optimization problem (10a), for which the gradient is defined as follows:
$\nabla C_{\lambda}\bigl{(}K\bigr{)}=\sum_{i\in\mathcal{V}}\nabla C_{\lambda,i}\bigl{(}K\bigr{)},$ (11)
where,
$\nabla C_{\lambda,i}\bigl{(}K\bigr{)}=\lambda\big{(}K-\widehat{K}_{i}\bigl{(}K\bigr{)}\big{)},$ (12a)
$\text{with}\ \ \widehat{K}_{i}\bigl{(}K\bigr{)}$ $=\operatorname*{arg\,min}_{\breve{K}\in\mathcal{K}_{i}}C_{i}\bigl{(}\breve{K}\bigr{)}+\frac{\lambda}{2}\|\breve{K}-K\|^{2}.$ (12b)
@@
The computation of (12a) requires solving (12b). The gradient for the augmented inner cost function is given by $\nabla C_{i}\bigl{(}\breve{K}\bigr{)}+\lambda\big{(}\breve{K}-K\big{)}$, where, $\nabla C_{i}\bigl{(}\breve{K}\bigr{)}$ can be obtained using (7) or the zeroth-order approach [3, Algorithm 1], and $K$ is the current estimate of the meta feedback policy. Thus, this approach is equally amenable to model-based or model-free scenarios.
As the cost induced by an estimate of the meta-policy at different systems is independent of each other, a client-server model can be used for computation as individual clients need not share their data. Once the inner optimizer is obtained, a gradient descent step is performed locally for the outer problem, and a local estimate for the meta-policy is obtained. Once all the systems have an estimate of the meta-policy, these estimates are communicated to the server. The meta-policy at the server is updated using estimates from different systems and the previous estimate from the server. This iterative gradient descent is repeatedly performed until the number of outer iterations guarantees the required estimate accuracy.
@@ 2       Communicate the policy at the server, $K_{i}^{s,0}\leftarrow K^{s},\text{ for all }i\in\mathcal{V}$
6                   $\|\widehat{K}_{i}\bigl{(}K_{i}^{s,p}\bigr{)}-\bar{K}\|\leqslant\delta$, c.f. (12b)
@@ III Convergence analysis
This section formally states the associated assumptions and auxiliary results for the convergence analysis for Algorithm 1.
The policy gradient method we use as a baseline for designing our algorithm finds the gradient at the current iterate of the feedback policy by simulating the system under that policy from random initial states. The assumption presented below ensures that all the states are visited with a non-zero probability while exploring.
@@
(Bound on initial cost): Let $K^{0}\in\cap\mathcal{K}_{i}$ be the policy used to initialize Algorithm 1, then $C(K^{0})<\infty$.
@@
(Bounded heterogeneity): There exists a constant $\sigma_{C}>0$ such that,
$\frac{1}{V}\sum_{i\in\mathcal{V}}\|\nabla C_{i}\bigl{(}K\bigr{)}-\nabla C\bigl{(}K\bigr{)}\|^{2}\leqslant\sigma_{C}^{2},$
for all $K\in\cap\mathcal{K}^{i}$ such that $C\bigl{(}K\bigr{)}<\infty$.
The motivation for this assumption is similar to that of [20, Lemma 4], where the authors had established a similar bound on the diversity of gradients based on the diversity in the system and cost matrices. In our case, the diversity in system matrices is bounded as the uncertainties $a$ and $b$ are bounded. Therefore, Assumption 5 aligns with the established theory.
Algorithm 1 requires the estimation of an optimizer of the inner problem (12b) up to a desired accuracy. The LQR cost as a function of the feedback policy satisfies the PL-inequality [3, Lemma 3], i.e., let $K_{i}^{*}$ be the optimal policy for a system realization $i\in\mathcal{V}$, then for any stable policy $K$, it holds that
@@
for some constant $\mu>0$. Now, for our approach, we add a quadratic regularizer term to this gradient-dominated function. We have established in Lemma 6 that the regularized cost (10b) also satisfies the PL inequality. This means it is possible to solve the inner minimization problem to any specified accuracy $\delta$ using gradient descent (or Gauss-Newton or Natural policy gradient) iterations of order $\mathcal{O}\mathopen{}\mathclose{{}\left(\log{(1/\delta)}}\right)$ [3, Theorem 7, Theorem 9].
Next, we build a sequence of auxiliary lemmas that will be useful in analyzing Algorithm 1. First, we show the gradient dominance property of the Moreau regularized cost in (10b).
@@
We extend the proof of [22, Proposition 4.1] to functions that satisfy PL inequality. For the policy $\breve{K}$ we have,
@@
The following lemma shows how the cost function’s smoothness property is inherited from the Moreau regularized cost function.
@@
Inequality (15) follows from [3, Lemma 25], where it was shown that for $\Delta_{K}:=\min\mathopen{}\mathclose{{}\left(\frac{\sigma_{\min}\bigl{(}Q{i}\bigr{)}}{4C_{i}\bigl{(}K_{2}\bigr{)}\|B_{i}\|\bigl{(}\|B_{i}+B_{i}K_{2}\|+1\bigr{)}},\|K_{2}\|}\right)$, $L$ is a polynomial in $\frac{C_{i}\mathopen{}\mathclose{{}\left(K^{0}}\right)}{\sigma_{\min}\bigl{(}\mathbb{E}_{x^{0}_{i}\sim\mathcal{D}^{i}}{\mathopen{}\mathclose{{}\left[x^{0}_{i}{x^{0}_{i}}^{\top}}\right]}\bigr{)}\sigma_{\min}\mathopen{}\mathclose{{}\left(Q}\right)}$, $\mathbb{E}{\mathopen{}\mathclose{{}\left[\|x^{0}_{i}\|^{2}}\right]}$, $\|A_{i}\|$, $\|B_{i}\|$, $\|R\|$, and ${1}/{\sigma_{\min}\mathopen{}\mathclose{{}\left(R}\right)}$. The second inequality (16) follows from [16, Lemma 4.5], where it was shown that the smoothness properties get transferred to the Moreau envelope regularized cost function.
@@
([18, Lemma 2], Bounded Heterogeneity of Moreau Envelope costs): For some policy $K\in\mathbb{R}^{m\times n}$, such that $C\bigl{(}K\bigr{)}<\infty$ and $\lambda>2\sqrt{2}L$, it follows that
@@
[18, Lemma 5], Cumulative gradient drift): For algorithm 1, at an outer iteration $s$ and under the assumption that $\theta\leqslant{\beta}/({2L_{C}})$, the following holds:
@@
The following proposition shows that the Moreau regularized cost for a system realization $i\in\mathcal{V}$ decreases during the iterations of the inner loop.
@@
Next, we present a corollary that establishes that when the policy generated by local iterations for a system realization is used for another system realization, the LQR cost incurred will be finite.
@@
(Finite cost by local policies): Let the conditions in Proposition 10 satisfy. Then, consider the local estimate $K_{i}^{s,P}$ at a system realization $i\in\mathcal{V}$ for an outer iteration $s$ of Algorithm 1, it follows that $C_{j}\bigl{(}K_{i}^{s,P}\bigr{)}<\infty$ for all $j\in\mathcal{V}$.
@@
(Finite cost at all iterations): Let the conditions in Proposition 10 satisfy. Then, for any iterate $K^{s}$ generated by Algorithm 1, $C_{i}\bigl{(}K^{s}\bigr{)}<\infty$ for all $i\in\mathcal{V}$.
@@
Proposition 10, Corollary 11, and Corollary 12 are analogous to the stability results in [20, Theorem 1, Theorem 2]. Through these, we show that the cost incurred by policies generated inside Algorithm 1 remains finite for all available system realizations.
@@
Corollary 15 shows that as the required accuracy of the optimizer increases (i.e., $\epsilon$ decreases), there is a quadratic increase in the number of outer iterations required, and the required inner accuracy also increases as a square root of the $\epsilon$. Note that the convergence result in [20] is up to a non-vanishing constant ball around the optimal solution ([20, Theorem 3, Theorem 4]) while the precision of our method is arbitrary by the precision of the inner optimization problem.

### v2

@@ I Introduction
Meta-learning facilitates quick adaptation to new tasks by applying previously acquired knowledge to new challenges [10, 11]. In RL, this evolves into Meta-RL, which trains agents to efficiently adapt to unknown environments [12, 13, 14, 15, 16]. The Meta-RL framework aims to develop a meta-policy that allows for rapid learning across various tasks, fostering knowledge adaptation to new scenarios.
@@
The meta-learning problem for LQR policy design was studied in [20, 21, 22] and focuses on meta-learning based on the MAML framework. MAML for the LQR problem in a single system and multi-task setting was studied in [22], which was later extended to a multi-system and multi-task setting by [20]. In their setup, a finite set of observed tasks is used to design an LQR meta-policy that can effectively and quickly adapt to unobserved tasks, but only local convergence is shown. Later, the authors in [21] studied system heterogeneity and provided global convergence guarantees.
When applying policy gradient techniques, MAML and ME define different surrogate cost functions and access different computational oracles. MAML uses the estimation of the gradient and Hessian of the cost function, whereas ME uses an approximate solution of an inner optimization problem via a first-order oracle. The inner loop accuracy parameter appears in the convergence guarantee, thereby providing the user with explicit control over the quality of the solution. Complied with this theoretical intuition, recent studies have shown that ME allows for a better empirical performance than MAML in multi-task setups [17, 19, 18].
@@
The main contribution of this paper is a novel first-order meta-RL algorithm (MEMLQR) that computes LQR meta-policies that can be efficiently adapted for an unseen system realization. By integrating meta-learning principles, our approach improves policy adaptability, ensuring effective performance in various scenarios and quick convergence.
@@
    We propose a first-order iterative algorithm to optimize the defined augmented cost in a client-server setup.
@@
    We show the convergence of the proposed algorithm to an approximate first-order stationary point. We also show that the policies generated by the algorithm will incur a finite cost for all the available system realizations.
@@ II-A Preliminaries
where, $\mathcal{A}:=\mathcal{A}_{0}+\sum_{i=1}^{p}a_{i}\mathcal{A}_{i},\mathcal{B}:=\mathcal{B}_{0}+\sum_{j=1}^{q}b_{j}\mathcal{B}_{j}$, $x^{t}\in\mathbb{R}^{n}$ is the state and $u^{t}\in\mathbb{R}^{m}$ is the input to the system at time $t$, $a:=\mathopen{}\mathclose{{}\left(a_{1}\;a_{2}\;\cdots\;a_{p}}\right)\in\mathbb{R}^{p}$ and $b:=\mathopen{}\mathclose{{}\left(b_{1}\;b_{2}\;\cdots\;b_{q}}\right)\in\mathbb{R}^{q}$ are bounded uncertainties that determine the system dynamics along with matrices $\mathcal{A}_{0}\in\mathbb{R}^{n\times n}$ and $\mathcal{B}_{0}\in\mathbb{R}^{n\times m}$, and sets of matrices, $\bar{\mathcal{A}}{}:=\mathopen{}\mathclose{{}\left(\mathcal{A}_{1},\mathcal{A}_{2},\cdots,\mathcal{A}_{p}}\right)\in\mathopen{}\mathclose{{}\left(\mathbb{R}^{n\times n}}\right)^{p}$ and ¯$\bar{\mathcal{B}}{}:=\mathopen{}\mathclose{{}\left(\mathcal{B}_{1},\mathcal{B}_{2},\cdots,\mathcal{B}_{q}}\right)\in\mathopen{}\mathclose{{}\left(\mathbb{R}^{n\times m}}\right)^{q}$. It is assumed that all possible realizations of the $\mathcal{A}-\mathcal{B}$ pair are controllable. Note that we adopt this model to be true to the earlier literature for uncertain linear systems [23]. Still, we emphasize the fact that a simpler model with uncertain system matrices (such as a norm bound around given matrices) may also be considered.
@@
subject to (2), where $Q\in\mathbb{R}^{n\times n}$ and $R\in\mathbb{R}^{m\times m}$ are the cost matrices such that $Q\succeq 0$ and $R\succ 0$. Without loss of generality, it is assumed that the cost matrices $Q$ and $R$ are the same for all the systems. This models a scenario where the controllers have the same goal or are trying to accomplish a similar task. The optimal control policy for the above cost function is a linear feedback policy of form $u^{t}_{i}=-K_{i}x^{t}_{i}$ for a policy $K_{i}\in\mathbb{R}^{m\times n}$ [1, Chapter 3]. Therefore, the cost can be expressed as
@@ II-B Moreau Envelope-based LQR Meta-Policy Estimation
While the approach presented by the policy gradient iterations in (5) is efficient for the computation of approximate optimal policies, here we focus on the task of learning good initialization that can be efficiently adapted for the LQR problem on unseen realizations of (2). We propose to use the Moreau Envelope (ME) approach for meta-learning [19, 18, 17] that defines an optimization problem
@@
We propose a gradient-based iterative framework termed MEMLQR: Moreau Envelope based Meta Linear Quadratic Regulator, shown in Algorithm 1, to solve the bilevel optimization problem (9a), for which the gradient is
$\hskip 48.36967pt\nabla C_{\lambda}\bigl{(}K\bigr{)}=\sum_{i\in\mathcal{V}}\nabla C_{\lambda,i}\bigl{(}K\bigr{)},$ (10a)
[…]
$\text{with}\ \ \widehat{K}_{i}\bigl{(}K\bigr{)}=\operatorname*{arg\,min}_{\breve{K}\in\mathcal{K}_{i}}C_{i}\bigl{(}\breve{K}\bigr{)}+\frac{\lambda}{2}\|\breve{K}-K\|^{2}.$ (10c)
@@
The computation of (10b) requires solving (10c). The gradient for the augmented inner cost function is given by $\nabla C_{i}\bigl{(}\breve{K}\bigr{)}+\lambda\big{(}\breve{K}-K\big{)}$, where, $\nabla C_{i}\bigl{(}\breve{K}\bigr{)}$ can be obtained using (6) or the zeroth-order approach [3, Algorithm 1], and $K$ is the current estimate of the meta feedback policy. Thus, this approach is equally amenable to model-based or model-free scenarios.
As the cost induced by an estimate of the meta-policy at different systems is independent of each other, a client-server model can be used for computation as individual clients need not share their data. Once the inner optimizer is obtained, a gradient descent step is performed locally for the outer problem, and a local estimate for the meta-policy is obtained. Once all the systems have an estimate of the meta-policy, these estimates are communicated to the server. The meta-policy at the server is updated using estimates from different systems and the previous estimate from the server. This iterative gradient descent is performed until the number of outer iterations guarantees the required estimate accuracy.
@@ 2       Communicate the policy at the server, $K_{i}^{s,0}\leftarrow K^{s},\text{ for all }i\in\mathcal{V}$
6                   $\|\widehat{K}_{i}\bigl{(}K_{i}^{s,p}\bigr{)}-\bar{K}\|\leqslant\delta$, c.f. (10c)
@@ III Convergence analysis
This section formally states the associated assumptions and auxiliary results for the convergence analysis for Algorithm 1. Due to space considerations, the proofs have been moved to [24]. We show in our main result Theorem 14 that, after a threshold number of iterations, the algorithm converges to a neighbourhood of a first-order stationary point of the Moreau regularized cost $C_{\lambda}\bigl{(}\cdot\bigr{)}$. This neighbourhood is determined by the number of iterations, the regularizer, the system and cost matrices, and an error incurred in the inner loop calculations. The policy gradient method we use as a baseline for designing our algorithm finds the gradient at the current iterate of the feedback policy by simulating the system under that policy from random initial states. The assumption presented below ensures that all the states are visited with a non-zero probability while exploring.
@@
(Bound on initial cost): Let the set $\mathcal{K}:=\cap\mathcal{K}_{i}$ be non-empty and $K^{0}\in\mathcal{K}$ be the policy used to initialize Algorithm 1, then $C(K^{0})<\infty$.
@@
(Bounded heterogeneity): For all $K\in\mathcal{K}$ such that $C\bigl{(}K\bigr{)}<\infty$, there exists a constant $\sigma_{C}>0$ satisfying,
$\frac{1}{V}\sum_{i\in\mathcal{V}}\|\nabla C_{i}\bigl{(}K\bigr{)}-\nabla C\bigl{(}K\bigr{)}\|^{2}\leqslant\sigma_{C}^{2}.$
The motivation for this assumption is similar to that of [21, Lemma 4], where the authors had established a similar bound on the diversity of gradients based on the diversity in the system and cost matrices. In our case, the diversity in system matrices is bounded as the uncertainties $a$ and $b$ are bounded. Therefore, Assumption 5 aligns with the established theory. Algorithm 1 requires the estimation of an optimizer of the inner problem (10c) up to a desired accuracy. The LQR cost as a function of the feedback policy satisfies the PL-inequality [3, Lemma 3], i.e., let $K_{i}^{*}$ be the optimal policy for a system realization $i\in\mathcal{V}$, then for any stable policy $K$ and some constant $\mu>0$, it holds that
@@
In our approach, we add a quadratic regularizer term to this gradient-dominated function. We have established in Lemma 6 that the regularized cost (9b) also satisfies the PL inequality. This means it is possible to solve the inner minimization problem to any specified accuracy $\delta$ using gradient descent (or Gauss-Newton or Natural policy gradient) iterations of order $\mathcal{O}\mathopen{}\mathclose{{}\left(\log{(1/\delta)}}\right)$ [3, Theorem 7, Theorem 9]. Next, we build a sequence of auxiliary lemmas that will be useful in analyzing Algorithm 1. First, we show the gradient dominance property of the Moreau regularized cost in (9b).
@@
We extend the proof of [25, Proposition 4.1] to functions that satisfy PL inequality. For the policy $\breve{K}$ we have,
@@
The following lemma shows how the cost function’s smoothness property is inherited from the Moreau regularized cost.
@@
Inequality (13) follows from [3, Lemma 25], where it was shown that for $\Delta_{K}:=\min\mathopen{}\mathclose{{}\left(\frac{\sigma_{\min}\bigl{(}Q{i}\bigr{)}}{4C_{i}\bigl{(}K_{2}\bigr{)}\|B_{i}\|\bigl{(}\|B_{i}+B_{i}K_{2}\|+1\bigr{)}},\|K_{2}\|}\right)$, $L$ is a polynomial in $\frac{C_{i}\mathopen{}\mathclose{{}\left(K^{0}}\right)}{\sigma_{\min}\bigl{(}\mathbb{E}_{x^{0}_{i}\sim\mathcal{D}^{i}}{\mathopen{}\mathclose{{}\left[x^{0}_{i}{x^{0}_{i}}^{\top}}\right]}\bigr{)}\sigma_{\min}\mathopen{}\mathclose{{}\left(Q}\right)}$, $\mathbb{E}{\mathopen{}\mathclose{{}\left[\|x^{0}_{i}\|^{2}}\right]}$, $\|A_{i}\|$, $\|B_{i}\|$, $\|R\|$, and ${1}/{\sigma_{\min}\mathopen{}\mathclose{{}\left(R}\right)}$. The second inequality (14) follows from [17, Lemma 4.5], where it was shown that the smoothness properties get transferred to the Moreau envelope regularized cost function.
@@
([19, Lemma 2], Bounded Heterogeneity of Moreau Envelope costs): For some policy $K\in\mathbb{R}^{m\times n}$, such that $C\bigl{(}K\bigr{)}<\infty$ and $\lambda>2\sqrt{2}L$, it follows that
@@
[19, Lemma 5], Cumulative gradient drift): For algorithm 1, at an outer iteration $s$ and under the assumption that $\theta\leqslant{\beta}/({2L_{C}})$, the following holds:
@@
The following proposition shows that Moreau regularized cost for a realization $i\in\mathcal{V}$ decreases during the iterations of the inner loop.
@@
Next, we present a corollary that establishes that the policy generated by local iterations for a system realization, when used for another system realization, incurs a finite cost.
@@
(Finite cost by local policies): Let the conditions in Proposition 10 be satisfied. Then, consider the local estimate $K_{i}^{s,P}$ at a system realization $i\in\mathcal{V}$ for an outer iteration $s$ of Algorithm 1, it follows that $C_{j}\bigl{(}K_{i}^{s,P}\bigr{)}<\infty$ for all $j\in\mathcal{V}$.
@@
(Finite cost at all iterations): Let the conditions in Proposition 10 be satisfied. Then, for any iterate $K^{s}$ generated by Algorithm 1, $C_{i}\bigl{(}K^{s}\bigr{)}<\infty$ for all $i\in\mathcal{V}$.
@@
Proposition 10, Corollary 11, and Corollary 12 are analogous to the stability results in [21, Theorem 1, Theorem 2]. Through these, we show that the cost incurred by policies generated inside Algorithm 1 remains finite for all available system realizations.
@@
Corollary 15 shows that as the required accuracy increases (i.e., $\epsilon$ decreases), there is a quadratic increase in the number of outer iterations required, and the required inner accuracy also increases as $\sqrt{\epsilon}$. Note that the convergence result in [21] is up to a non-vanishing constant ball around the optimal solution ([21, Theorem 3, Theorem 4]) while the precision of our method is arbitrary by the precision of the inner optimization problem.